Which of the following is a good way to quickly determine whether 3⁄12 and ¼ are equal?

Cross multiplication

Inverse multiplication

Reciprocal multiplication

Equal multiplication
Flag Question
1
Expert's answer
2019-08-29T08:51:35-0400

A good way to quickly determine whether 3⁄12 and ¼ are equal will be


Cross multiplication


"\frac { 3} {12 }= \frac { 1} {4 } \Leftrightarrow 3*4=12*1 \Leftrightarrow 12=12"

Answer: Cross multiplication.